OpenStudy (anonymous):

a^2 -1a-6=0.....what are the factors here?

OpenStudy (saifoo.khan):

3 or-2

OpenStudy (anonymous):

factors are \[a^2-a-6=(a-3)(a+2)=0\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw